OpenStudy (anonymous):

A single tree produces about 2.6 × 10^2 lb of oxygen each year. The Amazon rainforest has about 3.9 × 10^11 trees. About how many pounds of oxygen are produced in the rainforest each day ? 2.8 × 10 lb

OpenStudy (anonymous):

ok so what you want to do is multiply the amount a tree produces in a year by the number of trees in the rainforest and then divide by 365. That should be your solution

OpenStudy (anonymous):

2.6x10^2=260 3.9x10^11=390000000000 260x390000000000=101400000000000 101400000000000/365=277808219178.0822 so 277808219178.0822 is my answer

OpenStudy (anonymous):

it's not my answer

OpenStudy (anonymous):

bacause of this 2.8 × 10 __ lb

OpenStudy (anonymous):

it may be looking for the answer in scientific notation, try 2.8*10^11

OpenStudy (anonymous):

ok, yes it wants scientific notation. so 2.8 x 10^11

OpenStudy (anonymous):

Thank you